Find the equation of the line with the given points.
Answer y =
(1,6, - 3) and (5/6, 3)
Ok I get very confused with frractions and know that I am suppose to do the
3 - (-3)
5/6 - 1)
I am unsure first how to get that slope with the fraction there. Once I get the slope and put it in the correct form of y = mx + b, how would I get the b point?

Answer by Alan3354(69443)

Find the equation of the line with the given points.
Answer y =
(1/6,-3) and (5/6,3) ?? Is this what you meant?
Ok I get very confused with frractions and know that I am suppose to do the
3 - (-3) = 6 = diffy
5/6 - 1/6 = 2/3 = diffx
----------
slope m = diffy/diffx = 6/(2/3) = 9
Use y = mx + b and either point to find b
-3 = 9*(1/6) + b
b = -9/2
--------
y = 9x - 9/2